OpenStudy (anonymous):

the variables x and y vary directly. write an equation that relates x and y. then find x when y=-4. x=5, y=-15

OpenStudy (anonymous):

@saifoo.khan could you help me again? this my last one.

OpenStudy (saifoo.khan):

I think you should try this one..

OpenStudy (saifoo.khan):

I will check your working.

OpenStudy (anonymous):

ok -15=x(5) -75x y=-300 ?

OpenStudy (anonymous):

first part second part

OpenStudy (saifoo.khan):

-15=k(5) k = -3

OpenStudy (saifoo.khan):

Next: y = k x y = -3(x)

OpenStudy (anonymous):

y=12?

OpenStudy (saifoo.khan):

Next: y = k x y = -3(x) -12 = -3(x) Now solve for x.

OpenStudy (anonymous):

4

OpenStudy (saifoo.khan):

Right.

OpenStudy (anonymous):

is that it??

OpenStudy (saifoo.khan):

Yup.